OpenStudy (anonymous):

What are the solutions to the quadratic equation 4x^2 + 28x = 0?

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

the options x = 4 and x = −7 x = 4 and x = 7 x = 0 and x = −7 x = 0 and x = 7

Nnesha (nnesha):

would you like to factor it or do you want to use quadratic formula ?

OpenStudy (anonymous):

quadratic formula

Nnesha (nnesha):

alright then \[\huge\rm x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a}\] where a b c are \[\huge\rm Ax^2+Bx+C=0\] a= leading coefficient b=middle term c=constant term substitute

OpenStudy (anonymous):

so its, 4x^2 + 28x = 0? but theres no c

Nnesha (nnesha):

yes right so substitute c for 0

OpenStudy (anonymous):

Oh, okay

Nnesha (nnesha):

ye :D

OpenStudy (anonymous):

Would I plug in those numbers?

Nnesha (nnesha):

yes a= ?? b= ?

OpenStudy (anonymous):

a=4 and b=28, right

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

so then the whole equation is 4x^2 +28x +0=0

Nnesha (nnesha):

yes right so substitute a for 4 b for 28 and c for 0

OpenStudy (anonymous):

ye